19- Roman Destinations issues bonds due in 10 years with a stated interest rate of 8% and a face value of $410,000. Interest payments are made semi-annually. The market rate for this type of bond is 7%. Using present value tables, calculate the issue price of the bonds. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Multiple Choice

  • $410,000.

  • $224,831.

  • $439,137.

  • $383,330.
Answer #1
Number of periods 20 =10*2
Interest rate 3.50% =7%/2
Semi-annual interest 16400 =410000*8%/2
PV factor:
Semi-annual interest PVA of $1 14.2124 =(1-(1.035)^-20)/0.035
Principal PV of $1 0.50257 =1/1.035^20
Amount PV factor 3.5% Present value
Semi-annual interest 16400 14.2124 233083
Principal 410000 0.50257 206054
Total 439137
Option C $439,137 is correct